# A generalization of the Choi isomorphism with application to open quantum systems

### v1

@@ V.2 Comparison with Paulsen-Shultz-Kye-Han isomorphism
see (14). Hence an obvious generalization consists in replacing the orthonormal basis $\left(|i\rangle\langle j|\right)_{0\leq i,j<N}$ of ${\mathcal{A}}$ by a general one, say, $\left(F_{\alpha}\right)_{0\leq\alpha<N^{2}}$. This idea has been worked out in PS13 , K22 and HK23 ; therefore we will call the resulting isomorphism $\mbox{PSKH}({\mathcal{E}})$ the “Paulsen-Shultz-Kye-Han isomorphism" after the authors of these papers.
For $N=2$, the orthonormal basis $(F_{0},F_{1},F_{2},F_{3})$ defined in (61) and satisfying $F_{\alpha}^{\ast}=F_{\alpha}$, and the transposition $\phi$ we obtain for $\mbox{PSK}({\phi})$ the same result as for the dePillis-Jamiołski isomorphism (64). The Paulsen-Shultz-Kye-Han isomorphism hence provides a criterion for complete positivity that applies to some orthonormal bases, but not to all.
@@ V.3 Comparison with Frembs-Cavalcanti isomorphism
The linear space ${\mathcal{A}}$ equipped with the first product (66) can be identified with the $C^{\ast}$-algebra ${\mathcal{A}}$, the second product (67) defines a different $C^{\ast}$-algebra denoted by ${\mathcal{A}}^{\rm{op}}$. For both $C^{\ast}$-algebras the “star" is given by the adjoint map $\ast:{\mathcal{A}}\rightarrow{\mathcal{A}}$ which maps each operator onto its adjoint. There exist anti-linear isomorphisms between both $C^{\ast}$-algebras, for example, the adjoint map. With respect to a chosen orthonormal basis in ${\mathcal{H}}$ the transposition $\phi$ can be defined and represents a linear isomorphism $\phi:{\mathcal{A}}^{\rm{op}}\rightarrow{\mathcal{A}}$.
@@
$\phi\otimes{\mathbbm{1}}:{\mathcal{A}}^{\rm{op}}\otimes{\mathcal{A}}\rightarrow{\mathcal{A}}\otimes{\mathcal{A}}\;,$ (72)
where $\phi$ is the linear isomorphism $\phi:{\mathcal{A}}^{\rm{op}}\rightarrow{\mathcal{A}}$ given by the transposition of matrices. It follows that for $X\geq 0$ where $X\in{\mathcal{A}}^{\rm{op}}\otimes{\mathcal{A}}$ we have
@@
and hence $(\phi\otimes{\mathbbm{1}})(X)\geq 0$ in ${\mathcal{A}}\otimes{\mathcal{A}}$. The converse can be shown analogously.
Therefore $\rho\in{\mathcal{A}}^{\rm{op}}\otimes{\mathcal{A}}$ is positive iff $(\phi\otimes{\mathbbm{1}})(\rho)\in{\mathcal{A}}\otimes{\mathcal{A}}$ is positive. Inserting this result into Theorem 3 and using that $\phi\left(F_{\alpha}^{\ast}\right)=\overline{F_{\alpha}}$ we obtain the following:
@@
For simplicity, we will also call the isomorphism $\Phi\mapsto\rho(\Phi)$ the “Frembs-Cavalcanti isomorphism" and denote $\rho(\Phi)$ in (75) by $\mbox{FC}(\Phi)$. It is essentially the dePillis-Jamiołski isomorphism followed by the block transposition $\phi\otimes{\mathbbm{1}}$. In the Theorem 4 we got rid of the reference to the somewhat unfamiliar $C^{\ast}$-algebra ${\mathcal{A}}^{\rm{op}}$, but at the price of getting a matrix $\mbox{FC}(\Phi)$ instead of an operator.
For the example $N=2$ and $\Phi=\phi$ (transposition of matrices) we obtain for both orthonormal bases (59) and (61) the same matrix $\mbox{FC}(\phi)$ as for the Choi isomorphism (60). It is not positively semi-definite, in line with the fact that $\phi$ is not completely positive. On the other hand, $\mbox{FC}(\phi)$ is different from the GKS-matrix $G(\phi)$ given by (62), which shows that both isomorphisms, FC and GKS, are different.
@@ VI.1 Time-dependent GKS isomorphism
Motivated by the investigations of the dynamics of open quantum systems we consider a time dependent, trace-preserving completely positive map ${\mathcal{V}}(t)$ that operates on statistical operators $\rho_{0}\in{\sf B}^{+}_{1}({\mathcal{H}})$. The time-dependence is assumed to be sufficiently smooth. According to Theorem 2 we conclude
@@ VI.2 Time-dependent Lindblad equation
this implies that the remaining basis operators $\left(F_{\alpha}\right)_{1\leq\alpha<N^{2}}$ have vanishing trace. Splitting the double sum in (89) according to whether $\alpha,\beta=0$ or $>0$ we obtain
@@
$F(t)$ ::=:= $\sum_{1\leq\alpha<N^{2}}{\sf k}^{\alpha 0}(t)F_{\alpha}=:\widetilde{G(t)}-{\sf i}H(t),\;\widetilde{G(t)},H(t)\in{\sf B}({\mathcal{H}}),$ (92)
[…]
@@
since the diagonal matrix $\Delta(t)$ contains the given eigenvalues $\left({\sf k}^{\alpha}(t)\right)_{1\leq\alpha<N^{2}}$. The remaining entries of the full matrix ${\sf k}(t)\in{\sf B}({\mathbbm{C}}^{N^{2}})$ are reconstructed as follows. ${\sf k}^{00}(t)$ is uniquely determined by the condition Tr $\mbox{Tr }{\sf k}(t)=0$, see (82). From $H(t)$ and $G(t)$ we obtain $F(t)$ via (92) and (93). The expansion of $F(t)$ into the basis $\left(F_{\alpha}\right)_{1\leq\alpha<N^{2}}$ gives
@@ VI.3 Expansion up to $O(t^{2})$ terms
We make the general assumptions of Section VI.2 and hence consider a time evolution of the open quantum system given by eq. (88). The orthonormal basis $\left(F_{\alpha}\right)_{0\leq\alpha<N^{2}}$ of ${\mathcal{A}}$ is assumed to satisfy $F_{0}={\mathbbm{1}}/\sqrt{N}$ and, to further simplify the calculations, to consist of Hermitean operators, i. e., $F_{\alpha}\in{\sf B}_{0}({\mathcal{H}})$ for $1\leq\alpha<N^{2}$. Concerning the environment, we additionally assume that its Hilbert space is finite-dimensional, say, dim ($\mbox{dim }({\mathcal{H}}_{E})=M$. Moreover, we will restrict ourselves to the case where $\rho_{E}$ is a pure state. After introducing a suitable orthonormal base $\left(|i\rangle\right)_{0\leq i<M}$ in ${\mathcal{H}}_{E}$ we may write
@@
It will be convenient to also choose the associated orthonormal basis $\left(E_{\gamma}\right)_{0\leq\gamma\leq M^{2}}$ in $L({\mathcal{H}}_{E})$ defined analogously as in subsection IV.1. We will re-order this basis writing it in the way $\left(E_{\gamma}\right)_{-\lfloor M^{2}/2\rfloor\leq\gamma\leq\lfloor M^{2}/2\rfloor}$ for odd $M$ (for even $M$ one $\leq$ sign has to be replaced by $<$ in the following equations). Moreover, we choose
@@
$\mbox{Tr}_{E}\left(E_{\gamma}\,\rho_{E}\,E_{\delta}^{\ast}\right)$ $=$ $\mbox{Tr}_{E}\,|\gamma\rangle\langle 0|0\rangle\langle 0|0\rangle\langle\delta|=\delta_{\gamma\delta}&0\leq\gamma,\delta<M\\ 0&\mbox{else}\end{array}\right.\;,$ (109)
$\mbox{Tr}_{E}\left(E_{\delta}^{\ast}\,\rho_{E}\,E_{\gamma}\right)$ $=$ $\mbox{Tr}_{E}\,|\delta\rangle\langle 0|0\rangle\langle 0|0\rangle\langle\gamma|=\delta_{\gamma\delta}&-M<\gamma,\delta\leq 0\\ 0&\mbox{else}\end{array}\right.\;.$ (112)
@@
In what follows summations over $\alpha,\beta,\ldots$ will always run from $1$ to $N^{2}-1$ and over $\gamma,\delta,\ldots$ from $-\lfloor M^{2}/2\rfloor$ to $\lfloor M^{2}/2\rfloor$ if not indicated otherwise. Up to terms of order $O(t^{2})$ we have
@@
VI.3.1 00th order in $t$
First consider the 00th order. Here we have
@@
and hence the intial GKS-matrix $g_{0}^{\alpha,\beta}$ has the only non-vanishing entry $g_{0}^{00}=N$. Its eigenvalues are hence $N,0$, where the eigenvalue 00 is $(N^{2}-1)$-fold degenerate.
VI.3.2 $1$st order in $t$
@@
$-{\sf i}\,\mbox{Tr}_{E}\left[H_{S}\otimes{\mathbbm{1}}_{E},\rho_{0}\otimes\rho_{E}\right]=-{\sf i}\,\left[H_{S},\rho_{0}\right]=-{\sf i}\,\sum_{\alpha}h^{\alpha}\left(F_{\alpha}\,\rho_{0}-\rho_{0}\,F_{\alpha}\right)\;.$ (126)
$-{\sf i}\,\mbox{Tr}_{E}\left[{\mathbbm{1}}\otimes H_{E},\rho_{0}\otimes\rho_{E}\right]=-{\sf i}\,\mbox{Tr}_{E}\rho_{0}\otimes\left[H_{E},\rho_{E}\right]=-{\sf i}\,\rho_{0}\,\mbox{Tr}_{E}\left[H_{E},\rho_{E}\right]=0\;.$ (127)
@@
It follows that the $O(t)$ corrections $g_{1}^{\alpha,\beta}$ to the GKS-matrix have only non-zero contributions in the 00th row or 00th column, i. e., for $g_{1}^{0\alpha}$ and $g_{1}^{\alpha 0}$:
@@
VI.3.3 $2$nd order in $t$
For the second order in $t$ we will focus on the contributions to the GKS matrix for $\alpha,\beta\geq 1$. The contributions for $\alpha=0$ or $\beta=0$ will give third order corrections to the degenerate eigenvalue 00 and second order corrections to the dominant eigenvalue $N$. The latter does not jeopardize the positivity of the dominant eigenvalue as long as $t$ remains small. For details, we refer the reader to the following perturbation-theoretical calculations.
Proceeding analogously as in the case of linear order we would have to calculate nine terms which result from inserting the three terms of the Hamiltonian (114) into (123). Here we will confine ourselves to present only the non-vanishing results. The omission of terms that give only contributions to the 00th row or 00th column of $g_{2}$ will be indicated by the symbol $\stackrel{{\scriptstyle\alpha,\beta\geq 1}}{{\longrightarrow}}$. These terms, except the contribution to $g_{2}^{00}$, give corrections to the eigenvalues of $g(t)$ of maximal order $O(t^{3})$ and hence can be dismissed. The $g_{2}^{00}$-term can be disregarded since it will only yield a $O(t^{2})$-correction to the dominant eigenvalue $N$ of $g(t)$ which hence remains positive.
@@ VI.3.3 $2$nd order in $t$
$-\mbox{Tr}_{E}\left[H_{S}\otimes{\mathbbm{1}}_{E},\left[H_{SE},\rho_{0}\otimes\rho_{E}\right]\right]\stackrel{{\scriptstyle\alpha,\beta\geq 1}}{{\longrightarrow}}\sum_{\alpha\beta}\left(v^{\alpha 0}\,h^{\beta}+v^{\beta 0}\,h^{\alpha}\right)F_{\alpha}\rho_{0}F_{\beta}\;,$ (142)

### v2

@@ V.2 Comparison with Paulsen-Shultz-Kye-Han isomorphism
see (14). Hence an obvious generalization consists in replacing the orthonormal basis $\left(|i\rangle\langle j|\right)_{0\leq i,j<N}$ of $\mathcal{A}$} by a general one, say, $\left(F_{\alpha}\right)_{0\leq\alpha<N^{2}}$. This idea has been worked out in PS13 , therefore we will call the resulting isomorphism
$\mbox{PS}:L\left(L\left({\mathbbm{C}}^{N}\right),L\left({\mathbbm{C}}^{M}\right)\right)\rightarrow\left({\mathbbm{C}}^{N}\right)\otimes\left({\mathbbm{C}}^{M}\right)$ (67)
the “Paulsen-Shultz isomorphism" after the authors of this paper.
For $N=M=2$, the orthonormal basis $(F_{0},F_{1},F_{2},F_{3})$ defined in (62) and satisfying $F_{\alpha}^{\ast}=F_{\alpha}$, and any linear map ${\mathcal{E}}:L\left({\mathbbm{C}}^{2}\right)\rightarrow L\left({\mathbbm{C}}^{2}\right)$ we obtain for $\mbox{PS}({\mathcal{E}})$ the same result as for the dePillis-Jamiołski isomorphism (65). The Paulsen-Shultz isomorphism hence provides a criterion for complete positivity that applies to some orthonormal bases, but not to all. In fact, Paulsen and Shultz provide a condition for a general basis such that the Paulsen-Shultz isomorphism provides a criterion for complete positivity analogous to the the Choi criterion, see corollary 12 in PS13 , first item.
A further generalization is achieved in K22 ,K22 and HK24 by considering the matrix $C_{\phi}^{\sigma}$ defined by
$C_{\phi}^{\sigma}=\sum_{ij}e_{i,j}\otimes\phi(\sigma(e_{i,j}))\;,$ (68)
where $e_{i,j}=|i\rangle\langle j|$ denotes the standard matrix units and $\sigma\in L({\mathcal{A}})$ is given by $\sigma(A)=sAs^{\ast}$ for some invertible operator $s\in{\mathcal{A}}$. Here, too, the dependence on the special basis $e_{i,j}$ can be overcome since $C_{\phi}^{\sigma}$ depends only on the bilinear form $\langle,\rangle$, which is given by $\langle e_{i,j},\sigma(e_{k,l})\rangle=\delta_{ik}\delta_{jl}$. The linear transformation $\phi\mapsto C_{\phi}^{\sigma}$ will be called the “Paulsen-Shultz-Kye-Han isomorphism". It fulfills a refined and graded criterion for the k-positivity of $\phi$, see theorem 4.2 in HK23 .
Next we will consider the relation to the GKS-isomorphism. Let $U\in{\sf U}\left({\mathbb{C}}^{N}\right)$ be a unitary $N\times N$ matrix and $g\in L({\mathbb{C}}^{N\times N})$ the GKS-matrix of $\phi\in L({\mathcal{A}})$ w. r. t. the orthonormal basis $\left(E_{\alpha}U\right)_{0\leq\alpha<{N\times N}}$ in $\mathcal{A}$} and $E_{\alpha}$ defined in (20). It follows that
$\phi(U^{\ast}AU)$ $=$ $\sum_{\alpha\beta}g^{\alpha\beta}\,\left(E_{\alpha}U\right)\,\left(U^{\ast}AU\right)\,\left(E_{\beta}U\right)^{\ast}$ (69)
$=$ $\sum_{\alpha\beta}g^{\alpha\beta}\,E_{\alpha}\,A\,E_{\beta}^{\ast}\;.$ (70)
Hence the GKS-matrix $g$ coincides with the Choi matrix $C_{\phi}^{\sigma}$ in HK23 for $\sigma(A)=U^{\ast}AU$.
Summarizing, we conclude that for the special choice of orthonormal bases of the form $\left(E_{\alpha}U\right)_{0\leq\alpha<N}$ the GKS-isomorphism and the PSKH-isomorphism coincide, but in general they represent different generalizations of the Choi matrix isomorphism.
@@ V.3 Comparison with Frembs-Cavalcanti isomorphism
The linear space $\mathcal{A}$} equipped with the first product (71) can be identified with the $C^{\ast}$-algebra $\mathcal{A}$}, the second product (72) defines a different $C^{\ast}$-algebra denoted by ${\mathcal{A}}^{\rm{op}}$. For both $C^{\ast}$-algebras the “star" is given by the adjoint map $\ast:{\mathcal{A}}\rightarrow{\mathcal{A}}$ which maps each operator onto its adjoint. There exist anti-linear isomorphisms between both $C^{\ast}$-algebras, for example, the adjoint map. With respect to a chosen orthonormal basis in $\mathcal{H}$} the transposition $\tau$ can be defined and represents a linear isomorphism $\tau:{\mathcal{A}}^{\rm{op}}\rightarrow{\mathcal{A}}$.
@@
$\tau\otimes{\mathbbm{1}}:{\mathcal{A}}^{\rm{op}}\otimes{\mathcal{A}}\rightarrow{\mathcal{A}}\otimes{\mathcal{A}}\;,$ (77)
where $\tau$ is the linear isomorphism $\tau:{\mathcal{A}}^{\rm{op}}\rightarrow{\mathcal{A}}$ given by the transposition of matrices. It follows that for $X\geq 0$ where $X\in{\mathcal{A}}^{\rm{op}}\otimes{\mathcal{A}}$ we have
@@
and hence $(\tau\otimes{\mathbbm{1}})(X)\geq 0$ in ${\mathcal{A}}\otimes{\mathcal{A}}$. The converse can be shown analogously.
Therefore $\rho\in{\mathcal{A}}^{\rm{op}}\otimes{\mathcal{A}}$ is positive iff $(\tau\otimes{\mathbbm{1}})(\rho)\in{\mathcal{A}}\otimes{\mathcal{A}}$ is positive. Inserting this result into Theorem 3 and using that $\tau\left(F_{\alpha}^{\ast}\right)=\overline{F_{\alpha}}$ we obtain the following:
@@
For simplicity, we will also call the isomorphism $\Phi\mapsto\rho(\Phi)$ the “Frembs-Cavalcanti isomorphism" and denote $\rho(\Phi)$ in (80) by $\mbox{FC}(\Phi)$. It is essentially the dePillis-Jamiołski isomorphism followed by the block transposition $\tau\otimes{\mathbbm{1}}$. In the Theorem 4 we got rid of the reference to the somewhat unfamiliar $C^{\ast}$-algebra ${\mathcal{A}}^{\rm{op}}$, but at the price of getting a matrix $\mbox{FC}(\Phi)$ instead of an operator. The result of Theorem 4 is not new but can already be found in PS13 , theorem 22. In this sense the Frembs-Cavalcanti isomorphism is equivalent to a variant of the Poulsen-Shultz-isomorphism. For an earlier source, see also S01 , eq. (8).
For the example $N=2$ and $\Phi=\tau$ (transposition of matrices) we obtain for both orthonormal bases (60) and (62) the same matrix $\mbox{FC}(\tau)$ as for the Choi isomorphism (61). It is not positively semi-definite, in line with the fact that $\tau$ is not completely positive. On the other hand, $\mbox{FC}(\tau)$ is different from the GKS-matrix $G(\tau)$ given by (63), which shows that both isomorphisms, FC and GKS, are different.
@@ VI.1 Time-dependent GKS isomorphism
[…]
@@ VI.2 Time-dependent Lindblad equation
this implies that the remaining basis operators $\left(F_{\alpha}\right)_{1\leq\alpha<N^{2}}$ have vanishing trace. Splitting the double sum in (94) according to whether $\alpha,\beta=0$ or >0>0 we obtain
@@
$F(t)$ $:=$ $\sum_{1\leq\alpha<N^{2}}{\sf k}^{\alpha 0}(t)F_{\alpha}=:\widetilde{G(t)}-{\sf i}H(t),\;\widetilde{G(t)},H(t)\in{\sf B}({\mathcal{H}}),$ (97)
$G(t)$ $:=$ $\widetilde{G(t)}+{\sf k}^{00}(t)/(2N)\,{\mathbbm{1}}\;,$ (98)
@@
since the diagonal matrix $\Delta(t)$ contains the given eigenvalues $\left({\sf k}^{\alpha}(t)\right)_{1\leq\alpha<N^{2}}$. The remaining entries of the full matrix ${\sf k}(t)\in{\sf B}({\mathbbm{C}}^{N^{2}})$ are reconstructed as follows. ${\sf k}^{00}(t)$ is uniquely determined by the condition $\mbox{Tr }{\sf k}(t)=0$, see (87). From $H(t)$ and $G(t)$ we obtain $F(t)$ via (97) and (98). The expansion of $F(t)$ into the basis $\left(F_{\alpha}\right)_{1\leq\alpha<N^{2}}$ gives
@@ VI.3 Expansion up to $O(t^{2})$ terms
We make the general assumptions of Section VI.2 and hence consider a time evolution of the open quantum system given by eq. (93). The orthonormal basis $\left(F_{\alpha}\right)_{0\leq\alpha<N^{2}}$ of $\mathcal{A}$} is assumed to satisfy $F_{0}={\mathbbm{1}}/\sqrt{N}$ and, to further simplify the calculations, to consist of Hermitean operators, i. e., $F_{\alpha}\in{\sf B}_{0}({\mathcal{H}})$ for $1\leq\alpha<N^{2}$. Concerning the environment, we additionally assume that its Hilbert space is finite-dimensional, say, $\mbox{dim }({\mathcal{H}}_{E})=M$. Moreover, we will restrict ourselves to the case where $\rho_{E}$ is a pure state. After introducing a suitable orthonormal base $\left(|i\rangle\right)_{0\leq i<M}$ in ${\mathcal{H}}_{E}$ we may write
@@
It will be convenient to also choose the associated orthonormal basis $\left(E_{\gamma}\right)_{0\leq\gamma\leq M^{2}}$ in $L({\mathcal{H}}_{E})$ defined analogously as in subsection IV.1. We will re-order this basis writing it in the way $\left(E_{\gamma}\right)_{-\lfloor M^{2}/2\rfloor\leq\gamma\leq\lfloor M^{2}/2\rfloor}$ for odd $M$ (for even $M$ one $\leq$ sign has to be replaced by << in the following equations). Moreover, we choose
@@
[…]
@@
In what follows summations over $\alpha,\beta,\ldots$ will always run from $1$ to $N^{2}-1$ and over $\gamma,\delta,\ldots$ from $\lfloor M^{2}/2\rfloor$ to $\lfloor M^{2}/2\rfloor$ if not indicated otherwise. Up to terms of order $O(t^{2})$ we have
@@
VI.3.1 0th order in $t$
First consider the 0th order. Here we have
@@
and hence the initial GKS-matrix $g_{0}^{\alpha,\beta}$ has the only non-vanishing entry $g_{0}^{00}=N$. Its eigenvalues are hence $N,0$, where the eigenvalue 0 is $(N^{2}-1)$-fold degenerate.
VI.3.2 11st order in $t$
@@
${\sf i}\,\mbox{Tr}_{E}\left[H_{S}\otimes{\mathbbm{1}}_{E},\rho_{0}\otimes\rho_{E}\right]=-{\sf i}\,\left[H_{S},\rho_{0}\right]=-{\sf i}\,\sum_{\alpha}h^{\alpha}\left(F_{\alpha}\,\rho_{0}-\rho_{0}\,F_{\alpha}\right)\;.$ (131)
${\sf i}\,\mbox{Tr}_{E}\left[{\mathbbm{1}}\otimes H_{E},\rho_{0}\otimes\rho_{E}\right]=-{\sf i}\,\mbox{Tr}_{E}\rho_{0}\otimes\left[H_{E},\rho_{E}\right]=-{\sf i}\,\rho_{0}\,\mbox{Tr}_{E}\left[H_{E},\rho_{E}\right]=0\;.$ (132)
@@
It follows that the $O(t)$ corrections $g_{1}^{\alpha,\beta}$ to the GKS-matrix have only non-zero contributions in the 0th row or 0th column, i. e., for $g_{1}^{0\alpha}$ and $g_{1}^{\alpha 0}$:
@@
VI.3.3 22nd order in $t$
For the second order in $t$ we will focus on the contributions to the GKS matrix for $\alpha,\beta\geq 1$. The contributions for $\alpha=0$ or $\beta=0$ will give third order corrections to the degenerate eigenvalue 0 and second order corrections to the dominant eigenvalue $N$. The latter does not jeopardize the positivity of the dominant eigenvalue as long as $t$ remains small. For details, we refer the reader to the following perturbation-theoretical calculations.
Proceeding analogously as in the case of linear order we would have to calculate nine terms which result from inserting the three terms of the Hamiltonian (119) into (128). Here we will confine ourselves to present only the non-vanishing results. The omission of terms that give only contributions to the 0th row or 0th column of $g_{2}$ will be indicated by the symbol $\stackrel{{\scriptstyle\alpha,\beta\geq 1}}{{\longrightarrow}}$. These terms, except the contribution to $g_{2}^{00}$, give corrections to the eigenvalues of $g(t)$ of maximal order $O(t^{3})$ and hence can be dismissed. The $g_{2}^{00}$-term can be disregarded since it will only yield a $O(t^{2})$-correction to the dominant eigenvalue $N$ of $g(t)$ which hence remains positive.
@@
$\mbox{Tr}_{E}\left[H_{S}\otimes{\mathbbm{1}}_{E},\left[H_{SE},\rho_{0}\otimes\rho_{E}\right]\right]\stackrel{{\scriptstyle\alpha,\beta\geq 1}}{{\longrightarrow}}\sum_{\alpha\beta}\left(v^{\alpha 0}\,h^{\beta}+v^{\beta 0}\,h^{\alpha}\right)F_{\alpha}\rho_{0}F_{\beta}\;,$ (147)